# Quantum-Assisted Vehicle Routing: Realizing QAOA-based Approach on Gate-Based Quantum Computer

### v2

@@ 5. Conclusion
Quantum computing is an emerging and exciting avenue of research that is now becoming available to tackle problems from a different perspective. In this work, we explored the application of QAOA to the VRP by formulating it as a QUBO problem by encoding constraints such as flow conservation and subtour elimination into the cost function. Our study examined the feasibility of executing QAOA-based VRP solutions on current quantum hardware, with a particular focus on the impact of problem size on circuit complexity and noise resilience.
Starting with a three-node VRP, the QUBO was cast as an Ising Hamiltonian for QAOA with depth 2. The variational parameters were optimized using a classical optimizer by executing the circuit on the quantum computer multiple times, and the converged values were used to sample the outcome of the final circuit. The most probable observed state corresponded with the correct solution to the problem indicating a successful result for QAOA.
However, for larger instances of the problem, it was found that while QAOA can, in principle, provide approximate solutions to VRP, the rapid growth in circuit depth, specifically two-qubit gate depth, poses a significant practical challenge due to the limitations of noisy quantum processors. This suggests that the problem is dependent on hardware improvements and algorithmic optimizations, particularly in reducing the number of multi-qubit operations and mitigating noise effects. Our experiments show that coefficient normalization and carefully tuned penalty values are essential for improving feasibility in QAOA-based VRP. These heuristics helped recover feasible solutions more reliably, highlighting the importance of parameter design in practical quantum optimization. Future work will focus on developing more efficient QUBO formulations and encoding schemes, employing other hybrid approaches such as VQE and operator backpropagation, transfer learning of the variational parameters, and dynamic circuits.
As quantum hardware continues to advance, these improvements will be vital in determining whether QAOA and related quantum algorithms can provide a competitive advantage over classical optimization methods for real-world applications.

### v3

@@ 5. Conclusion
Quantum computing is an emerging and exciting avenue of research that is now becoming available to tackle problems from a different perspective. In this work, we explored the application of QAOA to the VRP by formulating it as a QUBO problem by encoding constraints such as flow conservation and subtour elimination into the cost function. Our study examined the feasibility of executing QAOA-based VRP solutions on current quantum hardware, with a particular focus on the impact of problem size and formulation on circuit complexity and noise resilience.
Starting with a three-node VRP, the QUBO was cast as an Ising Hamiltonian for QAOA with depth 2. The variational parameters were optimized using a classical optimizer by executing the circuit on the quantum computer multiple times, and the converged values were used to sample the outcome of the final circuit. The most probable observed state corresponded with the known correct solution to the problem indicating a successful result for QAOA.
However, for larger instances of the problem, it was found that while QAOA can, in principle, provide approximate solutions to VRP, the rapid growth in circuit depth, specifically two-qubit gate depth, poses a significant practical challenge due to the limitations of noisy quantum processors. This suggests that the problem is dependent on hardware improvements and algorithmic optimizations, particularly in reducing the number of multi-qubit operations and mitigating noise effects. Our experiments show that coefficient normalization and carefully tuned penalty values are essential for improving feasibility in QAOA-based VRP. These heuristics helped recover feasible solutions more reliably for larger problems, highlighting the importance of parameter design in practical quantum optimization. Future work will focus on developing more efficient QUBO formulations and encoding schemes, employing other hybrid approaches such as VQE and operator backpropagation, transfer learning of the variational parameters, and dynamic circuits.
The primary contribution of this work lies in its experimental, hardware-grounded analysis of QAOA for VRP. Rather than proposing a new algorithmic variant, we systematically expose the practical bottlenecks that arise when mapping constrained routing problems onto real quantum processors. These results clarify why many prior QAOA VRP approaches struggle to produce feasible solutions in practice, and they outline concrete methodological directions for advancing quantum optimization beyond simulation-only studies. As quantum hardware continues to advance, these improvements will be vital in determining whether QAOA and related quantum algorithms can provide a competitive advantage over classical optimization methods for real-world applications.